# Cooperative Nonlinear Guidance Strategies for Guaranteed Pursuit-Evasion

### v1

@@ I Introduction
Geometrical approaches have also found application in three-agent pursuit-evasion scenarios. For instance, in [13, 14], a method centered on line-of-sight guidance, a three-point guidance strategy, was explored. These three points were defined as the evader, the defender, and the pursuer. The approach demonstrated that if the defender remains aligned with the line-of-sight connecting the evader and the pursuer, the interception of the pursuer is assured before it nears the evader. In a related vein, a modified version of the command-to-line-of-sight guidance approach, incorporating optimal control theory and velocity error feedback, was introduced in [15]. Furthermore, in [16], the authors presented a guidance strategy based on a barrier Lyapunov function to protect the evader from a pursuer. It is worth noting that the use of the barrier function imposed restrictions on some engagement variables, such as the defender’s initial heading angle, thereby potentially limiting the target set in which the game could terminate. Motivated by these results, the focus of the current paper is analyzing and presenting a simple and intuitive geometry-based solution for guaranteed pursuit-evasion. The merits of this work can be succinctly summarized as follows:
@@ II Problem Formulation
We consider a cooperative defense problem involving nonholonomic agents, namely, a pursuer (P), an evader (E), and a defender (D). The pursuer aims to intercept the evader. In contrast, the defender’s objective is to neutralize the pursuer before it reaches the vicinity of the evader. Thus, the evader and the defender are allies, whereas the pursuer is the opposition. Such a scenario leads to an engagement, as shown in Figure 1. The agents evolve according to
@@
where $[x_{i},y_{i}]^{\top}\in\mathbb{R}^{2}$, $v_{i}\in\mathbb{R}_{+}$, and $\gamma_{i}\in(-\pi,\pi]$ denote the position, speed, and the heading angle of the $i$th agent, respectively, whereas $a_{i}$ is its steering control (lateral acceleration), which is assumed to be bounded. Thus, $\left|a_{i}\right|\leq a_{i}^{\max}\in\mathbb{R}_{+}$. This is unlike previous studies where the agents have simple dynamics, and their instantaneous heading angles were used to control them.
@@
where $v_{r_{\mathrm{EP}}},v_{r_{\mathrm{DP}}},v_{r_{\mathrm{DE}}},v_{\lambda_{\mathrm{EP}}},v_{\lambda_{\mathrm{DP}}}$, and $v_{\lambda_{\mathrm{DE}}}$, represent the components of relative velocities of the relevant agents along and perpendicular to the corresponding LOS of their respective engagements. The quantities $\delta_{k}$ in (II) denote the corresponding lead angles and are defined as $\delta_{\mathrm{PE}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{EP}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{PD}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{DP}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{ED}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}$, $\delta_{\mathrm{DE}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}.$ Note that (2a)–(2b) describe the equations of motion between the evader and the pursuer, while (2c)–(2d) represent the same for the defender-pursuer pair. On the other hand, the cooperative engagement between the evader-defender pair is described by (2e)–(2f).
@@ III Main Results
Consider the three-body engagement described using (II). If the defender maintains a positive closing speed with respect to the pursuer and attains an angle $\chi\in\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$, then the pursuer’s capture is guaranteed before the evader could be captured.
@@
$=$ $~{}\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}$
@@
which is affine with the steering controls of the evader and the defender, whereas that of the pursuer is an unknown quantity. In order to nullify $\beta$, we need to design $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$ such that the evader-defender team cooperatively maneuvers to place the defender on a fixed $\chi^{\star}$ in the interval $\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$ regardless of the initial three-body engagement geometry.
@@
for some $k_{1}\in\mathbb{N}$. Thereafter, to eliminate the reaching phase and achieve a global exact-time convergence, we consider the outer-layer sliding manifold as
@@
for some $k_{2}\in\mathbb{N}$, and $h(0)$ is designed to ensure $\bar{\mathcal{S}}(0)=0$, which can be obtained from (15) as $h(0)=-\underaccent{\bar}{\mathcal{S}}(0)$. The fundamental concept behind eliminating the reaching phase is to ensure global exact-time convergence for the sliding variables and achieve a complete sliding mode response. This, in consequence, bolsters robustness, as sliding mode controllers typically experience diminished robustness during their reaching phase. We are now ready to present the joint cooperative effort of the evader-defender alliance next.
@@
Consider the three-body engagement described using (II), the dynamics of the error variable, (12), and the dual-layer sliding manifold in (13) and (15). If the joint cooperative maneuver of the evader-defender team,
$\mathcal{U}=-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}$ (17)
is designed as
@@
for some $\mathcal{K}>\sup_{t\geq 0}\dfrac{a_{\mathrm{P}}^{\max}}{r_{\mathrm{DP}}}$, then the evader-defender team cooperatively maneuvers such that the defender converges to the desired angle $\chi^{\star}\in\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$, within a time $t^{\star}$ prescribed prior to the engagement, regardless of the three-body initial engagement geometry.
@@
$\dot{V}=$ $\mathrm{sign}(\bar{\mathcal{S}})\left[\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}\right.$
@@
Now substituting the above expression of $a_{\mathrm{E}}$ in the cost function (26) yields
$\mathcal{C}=\sqrt{\left(\dfrac{r_{\mathrm{DE}}}{\cos\delta_{\mathrm{ED}}}\left[\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}-\mathcal{U}\right]\right)^{2}+a_{\mathrm{D}}^{2}},$ (30)
which can also be written as
$\mathcal{C}=$ $\left[\left(\dfrac{r_{\mathrm{DE}}}{\cos\delta_{\mathrm{ED}}}\left[\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}+\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right.\right.\right.$
$\left.\left.\left.+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dot{g}(t)+\dot{h}(t)+\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}})\right]\right)^{2}+a_{\mathrm{D}}^{2}\right]^{\frac{1}{2}},$ (31)
after substituting for $\mathcal{U}$ from (18).
Computing the partial derivative of $\mathcal{C}$, obtained above, with respect to $a_{\mathrm{D}}$ and equating it to zero leads to the expression of the defender’s lateral acceleration in (28). Thereafter, one may use (29) to arrive at (27). It can be readily verified using higher-order derivative tests that the values thus obtained are indeed minima. This concludes the proof. ∎
@@ IV Simulations
(c) Steering controls of the evader and the defender.
@@
We now demonstrate the performance of the proposed nonlinear guidance strategy using simulation for various three-body engagement scenarios. We assume $v_{\mathrm{P}}=v_{\mathrm{D}}=200$ m/s, whereas $v_{\mathrm{E}}=100$ m/s. In each case, we place the evader at the origin of the inertial place at the beginning of the engagement. The initial positions of the agents in the trajectory plots are denoted using square markers, whereas an intercept is represented using an asterisk. In our simulations, the pursuer and the defender possess the same maneuverability capabilities, limited to ±$\pm 20$ g. On the other hand, the evader can only apply a maximum steering control of $10$ g in either direction. Here g$=9.81$ m/s2 is the acceleration due to gravity. Thus, the pursuer and the defender are evenly matched in terms of both speed and acceleration capabilities. The design parameters are chosen to be $t_{1}=3$ s, $t^{\star}=6$ s, $k_{1}=6$, $k_{2}=3$ and $\mathcal{K}=5$. Such choices of $t_{1}$ and $t^{\star}$ ensure that the sliding manifold converges to zero within 3 s, whereas the error vanishes within 6 s.
@@
Consider a scenario where the pursuer is using a guidance strategy that is a function of the pursuer-evader LOS rate ($\dot{\lambda}_{\mathrm{EP}}$), e.g., proportional-navigation guidance. For the pursuer, such a strategy may also be optimal in certain cases as it aims to arrive on the collision course with minimum effort. This situation is shown in Figure 3. The defender and pursuer are initially located $400$ m and $5000$ m radially apart from the evader, with respective LOS angles of $-45^{\circ}$ and $0^{\circ}$. The initial heading angles of the agents are $\gamma_{\mathrm{E}}=45^{\circ}$, $\gamma_{\mathrm{D}}=0^{\circ}$, and $\gamma_{\mathrm{P}}=180^{\circ}$. Under these settings, the initial value of the angle $\chi$ is $228^{\circ}$. It can be observed from Figure 2(a) that based on different values of the desired $\chi^{\star}$, the agents maneuver differently. However, the pursuer is always intercepted before it can capture the evader. Figure 2(b) depicts the sliding manifold and the error profiles of various desired angles $\chi^{\star}$, which evidences that $\underaccent{\bar}{\mathcal{S}}$ converges within 3 s and $\beta$ nullifies within 6 s, as expected. Therefore, regardless of the three-body initial engagement geometry, the evader-defender team maneuvers to ensure the pursuer’s capture. Figures 2(c) and 2(d) illustrate the various control efforts of the agents. It is important to note that the efforts of the evader and the defender (individually and jointly) have small magnitudes in the endgame.
@@
Since the proposed defense strategy is independent of the pursuer’s maneuver, the defender is able to capture the pursuer even if it uses a different guidance law. Figure 4 portrays various cases when the pursuer uses proportional-navigation guidance (PNG), pure pursuit guidance (PPG), and deviated pursuit guidance (DPG) to capture the evader. For a fixed angle, $\chi^{\star}=180^{\circ}$, the defender always intercepts the pursuer midway and safeguards the evader. In Figure 4, the initial conditions are kept the same as that in the previous case, except $\chi^{\star}=180^{\circ}$. We also observe similar behaviors of the sliding manifold, error, and the agents’ control efforts.
@@
We now consider cases when the evader is stationary, so the pursuer can head directly toward it with an optimal effort once any heading angle errors vanish. In Figure 5, the pursuer starts at a distance of 5000 m from the evader with a LOS angle of $2^{\circ}$. The defender, on the other hand, starts at three different positions in three different cases. With respect to the evader, the defender is at a relative separation of $400$ m, $1500$ m, and $2000$ m with LOS angles of $-45^{\circ}$, $-10^{\circ}$, and $10^{\circ}$, respectively. The defender aims to attain $\chi^{\star}=180^{\circ}$ in each case to arrive directly between the evader and the pursuer such that the latter’s capture can be guaranteed. We notice that even if the evader does not maneuver, the defender maneuvers accordingly to satisfy the specific geometrical conditions for a constant $\chi^{\star}$, thereby guaranteeing pursuit-evasion. On the other hand, the situation in Figure 6 assumes a fixed position of the defender ($r_{\mathrm{DE}}=400$ m, $\lambda_{\mathrm{DE}}=-45^{\circ}$), whereas the pursuer is located at different positions in different cases. The initial configurations of the pursuer relative to the evader are $r_{\mathrm{EP}}=5000$ m with LOS angles of $-5^{\circ},0^{\circ},$ and $2^{\circ}$, respectively. Once again, we observe that the fixed angle $\chi^{\star}$ is achieved by the defender in $t^{\star}=6$ s and maintained thereafter to ensure the pursuer’s capture. This further bolsters the arguments presented in this work that pursuit-evasion is guaranteed regardless of the three-body initial engagement geometry and the knowledge of the pursuer’s strategy.

### v2

@@ I Introduction
Geometrical approaches have also found application in three-agent pursuit-evasion scenarios. For instance, in [13, 14], a method centered on line-of-sight guidance, a three-point guidance strategy, was explored. These three points were defined as the evader, the defender, and the pursuer. The approach demonstrated that if the defender remains aligned with the line-of-sight connecting the evader and the pursuer, the interception of the pursuer is assured before it nears the evader. In a related vein, a modified version of the command-to-line-of-sight guidance approach, incorporating optimal control theory and velocity error feedback, was introduced in [15]. Furthermore, in [16], the authors presented a guidance strategy based on a barrier Lyapunov function to protect the evader from a pursuer. It is worth noting that the use of the barrier function imposed restrictions on some engagement variables, such as the defender’s initial heading angle, thereby potentially limiting the target set in which the game could terminate. Motivated by these results, the focus of the current paper is analyzing and presenting a simple and intuitive geometry-based solution for guaranteed pursuit evasion. The merits of this work can be succinctly summarized as follows:
@@
Note that even if the pursuer is sufficiently close to the evader, the evader and defender will cooperatively maneuver such that the defender can intercept the pursuer. Of course, there could be a few cases where the pursuer is too close to the evader, while the defender is too far away from them. In such cases, it is apparent that the defender may not be able to defend the evader. However, it is also worthwhile to note that such cases are impractical to consider from a three-body aircraft defense scenario because if the defender is too far away, then the engagement, for all practical purposes, is only for two agents.
@@ II Problem Formulation
We consider a cooperative defense problem involving nonholonomic agents, namely, a pursuer (P), an evader (E), and a defender (D). The pursuer aims to intercept the evader. In contrast, the defender’s objective is to neutralize the pursuer before it reaches the vicinity of the evader. Thus, the evader and the defender are allies, whereas the pursuer is the opposition. Such a scenario leads to an planar (2D) engagement, as shown in Figure 1. The agents evolve according to
@@
where $[x_{i},y_{i}]^{\top}\in\mathbb{R}^{2}$, $v_{i}\in\mathbb{R}_{+}$, and $\gamma_{i}\in(-\pi,\pi]$ denote the position, speed, and the heading angle of the $i$th agent, respectively, whereas $a_{i}$ is its steering control (lateral acceleration), which is assumed to be bounded. Thus, $\left|a_{i}\right|\leq a_{i}^{\max}\in\mathbb{R}_{+}$. This is unlike previous studies where the agents have simple dynamics, and their instantaneous heading angles were used to control them [17, 18].
@@
where $v_{r_{\mathrm{EP}}},v_{r_{\mathrm{DP}}},v_{r_{\mathrm{DE}}},v_{\lambda_{\mathrm{EP}}},v_{\lambda_{\mathrm{DP}}}$, and $v_{\lambda_{\mathrm{DE}}}$, represent the components of relative velocities of the relevant agents along and perpendicular to the corresponding LOS of their respective engagements. The quantities $\delta_{k}$ in (2) denote the corresponding lead angles and are defined as $\delta_{\mathrm{PE}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{EP}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{EP}}$, $\delta_{\mathrm{PD}}=\gamma_{\mathrm{P}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{DP}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DP}}$, $\delta_{\mathrm{ED}}=\gamma_{\mathrm{E}}-\lambda_{\mathrm{DE}}$, $\delta_{\mathrm{DE}}=\gamma_{\mathrm{D}}-\lambda_{\mathrm{DE}}.$ Note that (2a)–(2b) describe the equations of motion between the evader and the pursuer, while (2c)–(2d) represent the same for the defender-pursuer pair. On the other hand, the cooperative engagement between the evader-defender pair is described by (2e)–(2f).
@@ III Main Results
Consider the three-body engagement described using (2). If the defender maintains a positive closing speed with respect to the pursuer and attains an angle $\chi\in\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$, then the pursuer’s capture is guaranteed before the evader could be captured.
@@
$=$ $~{}-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}$
@@
which is affine with the steering controls of the evader and the defender, whereas that of the pursuer is an unknown quantity. The dynamics of $\beta$ can be written as
$\ddot{\beta}=$ $~{}\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}+\mathcal{U},$ (13)
where $\mathcal{U}$ is defined as the joint cooperative maneuver (or net control effort) given by
$\mathcal{U}=-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}.$ (14)
In order to nullify $\beta$, we need to design $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$ such that the evader-defender team cooperatively maneuvers to place the defender on a fixed $\chi^{\star}$ in the interval $\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$ regardless of the initial three-body engagement geometry.
@@
for some $k_{1}\in\mathbb{N}$. Note that the variable $t^{\star}$ denotes the time instant at which the error variable $\beta(t)$ nullifies to zero. As a consequence, the defender attains a predefined angle $\chi^{\star}\in\left[\pi/2,3\pi/2\right]$ within the same time $t^{\star}$. Thereafter, to eliminate the reaching phase and achieve a global exact-time convergence, we consider the outer-layer sliding manifold as
@@
for some $k_{2}\in\mathbb{N}$, and $h(0)$ is designed to ensure $\bar{\mathcal{S}}(0)=0$, which can be obtained from (17) as $h(0)=-\underaccent{\bar}{\mathcal{S}}(0)$. In (18), $t_{1}$ is the exact time instant from which onward the inner layer sliding manifolds become identically equal to zero. In other words, the sliding mode is enforced on the inner-layer sliding manifold at time $t=t_{1}$. The fundamental concept behind eliminating the reaching phase is to ensure global exact-time convergence for the sliding variables and achieve a complete sliding mode response. This, in consequence, bolsters robustness, as sliding mode controllers typically experience diminished robustness during their reaching phase. We are now ready to present the joint cooperative effort of the evader-defender alliance next.
@@
Consider the three-body engagement described using (2), the dynamics of the error variable, (13), and the dual-layer sliding manifold in (15) and (17). If the joint cooperative maneuver of the evader-defender team is designed as
@@
for some $\mathcal{K}>\sup_{t\geq 0}\left({a_{\mathrm{P}}^{\max}}/{r_{\mathrm{DP}}}\right)$, then the evader-defender team cooperatively maneuvers such that the defender converges to the desired angle $\chi^{\star}\in\left[{\pi}/{2},{3\pi}/{2}\right]$, within a time $t^{\star}$ prescribed prior to the engagement, regardless of the three-body initial engagement geometry.
@@
$\dot{V}=$ $\mathrm{sign}(\bar{\mathcal{S}})\left[-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}\right.$
@@
Now substituting the above expression of $a_{\mathrm{E}}$ in the cost function (27) yields which can also be written as $\mathcal{C}=\sqrt{\mathfrak{B}_{1}\left(\mathfrak{B}_{2}a_{\mathrm{D}}-\mathcal{U}\right)^{2}+a_{\mathrm{D}}^{2}}$, where $\mathfrak{B}_{1}=\dfrac{r_{\mathrm{DE}}}{\cos\delta_{\mathrm{ED}}}$ and $\mathfrak{B}_{2}=\dfrac{r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DE}}r_{\mathrm{DP}}}$. Computing the partial derivative of $\mathcal{C}$ with respect to $a_{\mathrm{D}}$ and equating it to zero leads to
$\dfrac{\partial\mathcal{C}}{\partial a_{\mathrm{D}}}=\dfrac{1}{2\mathcal{C}}\left[2\mathfrak{B}_{1}\left(\mathfrak{B}_{2}a_{\mathrm{D}}-\mathcal{U}\right)\mathfrak{B}_{2}+2a_{\mathrm{D}}\right]=0$
$\implies\mathfrak{B}_{1}\mathfrak{B}_{2}^{2}a_{\mathrm{D}}+a_{\mathrm{D}}=\mathfrak{B}_{1}\mathcal{U}\implies a_{\mathrm{D}}=\dfrac{\mathfrak{B}_{1}\mathcal{U}}{1+\mathfrak{B}_{1}\mathfrak{B}_{2}^{2}},$ (31)
which on substituting the value of $\mathcal{U}$ from (19) and $\mathfrak{B}_{1}$ and $\mathfrak{B}_{2}$ leads to the expression of the defender’s lateral acceleration in (29). Thereafter, one may use (30) to arrive at (28). It can be readily verified using higher-order derivative tests that the values thus obtained are indeed minima. By following a similar procedure, one can obtain the value of $a_{\mathrm{E}}$. This concludes the proof. ∎
@@ IV Simulations
(c) Steering controls $a_{\rm D}$ and $a_{\rm E}$.
(d) $\mathcal{U}$ and $a_{\rm P}$.
@@
(a) Trajectories.
(b) Sliding manifold and the error.
(c) Steering controls $a_{\rm D}$ and $a_{\rm E}$.
(d) $\mathcal{U}$ and $a_{\rm P}$.
Figure 7: The defender intercepts the pursuer with reduced maximum acceleration
@@
We now demonstrate the performance of the proposed nonlinear guidance strategy using simulation for various three-body engagement scenarios. We assume $v_{\mathrm{P}}=v_{\mathrm{D}}=200$ m/s, whereas $v_{\mathrm{E}}=100$ m/s. In each case, we place the evader at the origin of the inertial place at the beginning of the engagement. The initial positions of the agents in the trajectory plots are denoted using square markers, whereas an intercept is represented using an asterisk. In the following plots, the pursuer’s trajectories and inputs are shown in red, while those of the evader and defender are shown in blue and green, respectively. In our simulations, the pursuer and the defender possess the same maneuverability capabilities, limited to ±$\pm 20$ g. On the other hand, the evader can only apply a maximum steering control of $10$ g in either direction. Here g$=9.81$ m/s2 is the acceleration due to gravity. Thus, the pursuer and the defender are evenly matched in terms of both speed and acceleration capabilities. The design parameters are chosen to be $t_{1}=3$ s, $t^{\star}=6$ s, $k_{1}=6$, $k_{2}=3$ and $\mathcal{K}=5$. Such choices of $t_{1}$ and $t^{\star}$ ensure that the sliding manifold converges to zero within 3 s, whereas the error vanishes within 6 s.
@@
Consider a scenario where the pursuer is using a guidance strategy that is a function of the pursuer-evader LOS rate ($\dot{\lambda}_{\mathrm{EP}}$), for example, proportional-navigation guidance. For the pursuer, such a strategy may also be optimal in certain cases as it aims to arrive on the collision course with minimum effort. The guidance command for the pursuer is given by $a_{\rm P}=-\mathscr{N}\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\rm EP}$, where $\mathscr{N}=3$ is a constant. This situation is shown in Figure 3. The defender and pursuer are initially located $400$ m and $5000$ m radially apart from the evader, with respective LOS angles of $-45^{\circ}$ and $0^{\circ}$. The initial heading angles of the agents are $\gamma_{\mathrm{E}}=45^{\circ}$, $\gamma_{\mathrm{D}}=0^{\circ}$, and $\gamma_{\mathrm{P}}=180^{\circ}$. Under these settings, the initial value of the angle $\chi$ is $228^{\circ}$. It can be observed from Figure 3(a) that based on different values of the desired $\chi^{\star}$, the agents maneuver differently. However, the pursuer is always intercepted before it can capture the evader. Figure 3(b) depicts the sliding manifold and the error profiles of various desired angles $\chi^{\star}$, which evidences that $\underaccent{\bar}{\mathcal{S}}$ converges within 3 s and $\beta$ nullifies within 6 s, as expected. Therefore, regardless of the three-body initial engagement geometry, the evader-defender team maneuvers to ensure the pursuer’s capture. Figures 3(c) and 3(d) illustrate the various control efforts of the agents. It is important to note that the efforts of the evader and the defender (individually and jointly) have small magnitudes in the endgame.
@@
Since the proposed defense strategy is independent of the pursuer’s maneuver, the defender is able to capture the pursuer even if it uses a different guidance law. Figure 4 portrays various cases when the pursuer uses proportional-navigation guidance (PNG), pure pursuit guidance (PPG), and deviated pursuit guidance (DPG) to capture the evader. The guidance command for PNG, PPG, and DPG are chosen to be $a_{\rm P}=-\mathscr{N}\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\rm EP}$, $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}\right)$, and $a_{\rm P}=v_{\mathrm{P}}\dot{\lambda}_{\rm EP}-0.1\left(\gamma_{\mathrm{P}}-\lambda_{\mathrm{EP}}-\delta\right)$, respectively with $\mathscr{N}=3$ and $\delta=20^{\circ}$. For a fixed angle, $\chi^{\star}=180^{\circ}$, the defender always intercepts the pursuer midway and safeguards the evader. In Figure 4, the initial conditions are kept the same as that in the previous case, except $\chi^{\star}=180^{\circ}$. We also observe similar behaviors of the sliding manifold, error, and the agents’ control efforts.
@@
We now consider cases when the evader is stationary, so the pursuer can head directly toward it with an optimal effort once any heading angle errors vanish. In Figure 5, the pursuer starts at a distance of 5000 m from the evader with a LOS angle of $2^{\circ}$. The defender, on the other hand, starts at three different positions in three different cases. With respect to the evader, the defender is at a relative separation of $400$ m, $1500$ m, and $2000$ m with LOS angles of $-45^{\circ}$, $-10^{\circ}$, and $10^{\circ}$, respectively. The defender aims to attain $\chi^{\star}=180^{\circ}$ in each case to arrive directly between the evader and the pursuer such that the latter’s capture can be guaranteed. We notice that even if the evader does not maneuver, the defender maneuvers accordingly to satisfy the specific geometrical conditions for a constant $\chi^{\star}$, thereby guaranteeing pursuit-evasion. On the other hand, the situation in Figure 6 assumes a fixed position of the defender ($r_{\mathrm{DE}}=400$ m, $\lambda_{\mathrm{DE}}=-45^{\circ}$), whereas the pursuer is located at different positions in different cases. The initial configurations of the pursuer relative to the evader are $r_{\mathrm{EP}}=5000$ m with LOS angles of $-5^{\circ},0^{\circ},$ and $2^{\circ}$, respectively. Once again, we observe that the fixed angle $\chi^{\star}$ is achieved by the defender in $t^{\star}=6$ s and maintained thereafter to ensure the pursuer’s capture. We now demonstrate the performance of the proposed strategy when the maximum steering controls of the defender and pursuer are chosen to be $8$ g, whereas the maximum steering control input for the evader is chosen to be $4$ g. While the other engagement settings are kept the same as for $\chi^{\star}=180^{\circ}$, and we depict the performance under reduced maximum capability through Figure 7(d). It can be observed that the defender intercepts the pursuer by following a similar behavior as when maximum accelerations are higher. This further bolsters the arguments presented in this work that pursuit-evasion is guaranteed regardless of the three-body initial engagement geometry and the knowledge of the pursuer’s strategy.